A 15 kg box is pulled at an angle of 15* from the horizontal with a force of 55N In the coefficient of kinetic friction is 0.25. What is the normal force?


1
Expert's answer
2020-12-14T12:13:26-0500

Net force along the axis, parallel to inclined plane is:

"F + m g \sin \alpha - F_f", where "F = 55 N" and "F_f = \mu N" is static friction, "\alpha = 15^{\circ}"is the angle of inclination.

Assuming, we are pulling the box with no acceleration, "F + m g \sin \alpha - \mu N = 0", from where "N = \frac{F + m g \sin \alpha}{\mu} \approx 372.3 N".